Calculate the pH of a 2 M solution of H2SO4.
1
Expert's answer
2020-05-05T03:39:57-0400

"H_2SO_4->H^++HSO_4^-"

2 0 0

2-2=0 2 2

"HSO_4^-->H^++SO_4^{2-}"

2 2 0

2-x 2+x x

"Ka_2=1.1\u00d710^{-2}"

"\frac{x(2+x)}{2-x}=0.011"

"x^2+2.011x-0.022=0"

"x=[H^+]=0.011"

"pH=-log[H^+]=-log(0.011)=1.96"

"pH=1.96"